Dr. Salter wants to invest $50,000 to earn $4400 per year. He can invest in b-rated bonds paying 12% and CD's paying 4% per year. How much should be invested in each to realize exactly $4400 in intrest per year?
Answer by Earlsdon(6294)

Let x = the amount invested in b-rated bonds at 12% then ($50,000-x) = the amount invested in CDs at 4%
The annual interest earned by the b-rated bonds is 12% of x.
The annual interest earned by the CDs is 4% of ($50,000-x)
The sum of these two amounts = $4,400 per annum.
You can write the equation to solve for x after changing the percentages to their equivalent decimal values. (12% = 0.12 and 4% = 0.04)
0.12x + 0.04($50,000-x) = $4,400 Simplify and solve for x.
0.12x + $2,000 - 0.04x = $4,400 Combine like-terms.
0.08x + $2,000 = $4,400 Subtract $2,000 from both sides.
0.08x = $2,400 Divide both sides by 0.08
x = $30,000 This is the amount invested in the b-rated bonds at 12% per annum.
$50,000 - x = $50,000 - $30,000 = $20,000 This is the amount invested in CDs at 4%.